a bullet of mass m=3g, travelling at a velocity of v= 500 m/s, imbeds itself in the wooden block of a ballistic pendulum. if the wooden block has a mass of M= 0.3kg, to what height does the "bullet-block" combination rise


1
Expert's answer
2021-03-10T17:16:16-0500

"u=\frac{mv}{m+M}=\frac{0.003\cdot 500}{0.303}=4.95~\frac ms,"

"\frac{(m+M)u^2}{2}=(m+M)gh,"

"h=\frac{u^2}{2g}=\frac{4.95^2}{2\cdot 9.81}=1.25~m."